OpenStudy (anonymous):

If g(x)= x to the second power + 3, find g(4). A. 11 B. 19 C. 16 D. 8

OpenStudy (anonymous):

@WoahItsJian can you help?

jimthompson5910 (jim_thompson5910):

Replace every x with 4 \[\Large g(x) = x^2 + 3\] \[\Large g(x) = (x)^2 + 3\] \[\Large g(\color{red}{x}) = (\color{red}{x})^2 + 3\] \[\Large g(\color{red}{4}) = (\color{red}{4})^2 + 3\] \[\Large g(\color{red}{4}) = ??\] I'll let you compute the result

OpenStudy (anonymous):

Is it 19?

jimthompson5910 (jim_thompson5910):

yep it's 19